# A framework for the use of generative modelling in non-equilibrium statistical mechanics

### v4

@@ Acknowledgements
This paper more fully develops the theory outlined in a previous manuscript [1], and we are grateful to those listed there for their contributions to preliminary discussions on the topic. The account here is largely based on a series of talks given by the third named author at the Department of Philosophy of the University of Vienna, the Laufer Center for Physical and Quantitative Biology, the 2024 Conference on Complex Systems, and the informal discussions that followed. We thank the organisers and audiences of both. DARS acknowledges support from the Einstein Chair programme at the Graduate Centre of the City University of New York. KJF and DARS acknowledge support from the VERSES Research Lab.
@@ 1 Introduction
In statistics, generative models are joint probability distributions that model the relationships between different variables—most often, observations and their causes, or data and a set of labels. In the study of open and interacting systems, the use of statistical models—capable of describing such relationships—seems to be particularly apt. One approach to making such models, using variational (approximate Bayesian) inference [2, 3, 4], describes how a physical object reflects the properties of its environment and vice versa. In this paper, we argue that generative models provide an especially useful foundation in scientific modelling generally, and especially in the study of systems exhibiting adaptation, morphogenesis, and other non-equilibrium phenomena. More specifically, we argue that generative modelling—premised on the use of variational inference to model the causes of the signals that a coupled system is receiving [5]—is particularly apt for the practice of modelling the activity of modelling itself; for instance, in modelling sophisticated self-organising systems, where we write down a representation of the model that we assume is in play in prediction and action. We use the ensuing variational free energy principle to illustrate how implementing models which (i) formalise how one object keeps track of another, and (ii) codify these relationships, can increase the fidelity of our mathematical modelling.
@@
One of the points this paper will emphasise is that using this framework does not necessarily presuppose that a system actually performs inference on its environment. It is a truism that physical systems need not explicitly calculate their trajectories of motion, to be modelled as pursuing such trajectories. We need not assume that the particle itself is literally performing inference when we build models under our framework. What is at stake is an ‘as if’ description: we simply assume that there exists a quantity that varies systematically with the dynamics of the system; it so happens that surprisal (or, equivalently, variational free energy) is a good choice of quantity for physical reasons, and that the minimisation of this quantity is mathematically equivalent to inference, in the sense of being an estimator. The fact that it is our model of the system which does inference and not necessarily the system itself is both a boon and a bane. It enables us to model the system using the more tractable gradient of variational free energy and the mathematics of statistical inference, rather than studying the possibly highly non-linear coupling between the dynamics of the two systems. However, it is not necessarily a faithful representation of the system in the literalist sense. (See also [6].)
@@
The structure of the paper is as follows. We first present a conceptual introduction to the FEP—and the way it leverages the relationships encoded by a generative model to build a scientific model of the dynamics of physical objects using free energy gradients. This is possible by modelling the property of physical systems that they look as if they engage in modelling to an external observer. To sketch the argument: the FEP provides us with tools from dynamical systems and information theory, allowing us to model any coupled random dynamical systems as statistical estimators, i.e., as engaging in the statistical estimation of some quantity pertaining to the set of objects with which it interacts. Under the FEP, one object that is coupled to any other can thus be read as a representation of the coupled or estimated object, albeit of a deflationary, qualified sort (discussed in [7]; see also [8, 3]). More formally, the FEP applies to generative models that contain a partition of the states or paths of a joint system (a so called “particular” partition, i.e., into objects or particles). This partition induces a statistical boundary called a Markov blanket [9], which captures conditional independence relations amongst subsets of the system [2, 10]. The FEP says that, when such Markov blanketed subsets of the system exist, those subsystems track each others’ statistical properties (dynamically) across the blanket [11]. Essentially, the FEP is a modelling framework which captures the following fact: models of coupled particles or factorised systems can be written as parameter estimators, and estimating the parameters of a distribution is a form of Bayesian inference.
@@
We then argue that there is no ambiguity between model and modelled real-world system in this formalism. We can think of the FEP, metaphorically, as a map of the part of the territory that behaves as if it were a map—or looks as such to an external observer. Accounts such as [6] have encouraged practitioners of the FEP to be explicit about this point. Viewing the theory this way, not only is there no conflation between our FEP-theoretic model of the system and the system itself, but in fact we arrive at a novel modelling method for coupled random dynamical systems. In particular, we do not claim that the objects in the system themselves must literally be performing inference. Rather, we describe the dynamics as doing inference, in the sense of being a statistical estimator whose states parameterise a variational density.
@@
In summary we propose that the FEP provides a map of any possible map whatsoever—of, or held by—a physical system. (This includes the use of the FEP itself by a modeller.) We suggest that this leads to constraints on all possible maps, or models of physical things, arising from what it means to be a map or model of a physical process at all. In the fashion of Jaynes, who proposed that doing statistical mechanics is doing inference about the probabilities of microstates given knowledge of macrostates [12, 13], the FEP is a natural expression of statistical physics as a process of making models. More broadly we claim that this echoes the sentiments of Wittgenstein, providing a constraint on modelling in terms of logically necessary preconditions for sense-making and the parsing of sensory streams [7].
@@
For those already familiar with the FEP, this paper can be thought of as going back to basics, emphasising its foundation in inverting generative models (written as SDEs in generalised coordinates), using a path integral scheme for minimising variational free energy [14].
@@ 2.1 Fundamentals
Here we will describe the proposed free energy principle (also directing the reader to prior literature such as [11, 3, 4] for a more extended look). The FEP aims to describe the physical properties of a system in terms of dynamics on a manifold of probability distributions (i.e., as variational inference) [15], resting on the useful property that physical objects generally look as if they track or infer the statistics of things to which they are coupled [3, 16, 11]. Formulated most generally, the FEP says that the average path taken by an object that is sparsely coupled to its environment is characterised by a minimisation of a variational free energy functional.
We consider a particle (also referred to as a ‘system’) with states $\mu$ coupled to another system with states $\eta$ (henceforth, the ‘environment’) through a set of variables $B$ with values $b$, and a steady state density $p(\eta,b,\mu)$ such that $\mu$ is conditionally independent of $\eta$ given $B$. After Pearl (see [9]) we refer to $B$ as the Markov blanket of the system. Typically, $B$ will itself consist of two different kinds of states—sensory states, providing the system signals about its environment, and active states, providing interventions of the system on the environment. The coupling at play puts us in a non-equilibrium situation, where a system can be driven by fluxes of energy or matter from other systems in its environment [17].
The FEP begins by saying that if there exists a boundary in the form of a Markov blanket, then there exists a function $\sigma$ relating (the dynamics of) internal states to the environment across the boundary, and in particular, such that $\sigma(\hat{\mu}_{b})=\hat{\eta}_{b}$, where $\hat{\,\cdot\,}_{b}$ denotes the conditional expectation given the blanket state $b$ [16, Lemma 4.3]. Following that, the aim of the FEP is to understand what the generative model $p(\eta,b,\mu)$ says about the dynamics of the subsystem with states $\mu$, beyond just its marginal distribution. In particular: if we assume that $\hat{\eta}_{b}$ parametrises $p(\eta\mid b)$, then we have made $\mu$ into an estimator of the statistics of the environment.
@@
in terms of a gradient flow on the surprisal weighted by a particular matrix field [18, 19, 20, 21], having the form
@@
Moreover, the log-probability is able to describe the probability of all fluctuations of the system in terms of their distance from the most likely state, as well as physically-derived quantities [22, 17, 23], making this a meaningful equation. (See [24] where this is applied to reproduce certain aspects of stochastic thermodynamics.)
@@
Inference by parameter estimation implies the KL divergence between the posterior and the true density vanishes as the parameter varies. We will write
@@
with equality precisely when the KL term vanishes. We conclude that the minimisation of surprisal implies the vanishing of the divergence term. Let us apply the Laplace approximation [25] such that the expected state is approximately the mode of the density. It follows that the most probable state is the least surprising one. Hence, for physical reasons, we expect the surprisal to be minimised. From that reasoning, (2.1.1) has the equivalent form
@@
What we have done is used the fact that the dynamics of a (steady state) system minimise surprisal on average, and argued that whenever equipped with a Markov blanket they also minimise variational free energy; and hence, the dynamics of a system with a particular partition can always be written as a gradient flow on variational free energy. This gradient flow is what we identify as actually being the process (i.e., the performance) of inference. Explicit examples of this can be found in [24, 26, 11]. The particular partition becomes crucial here, since this is what we leverage in our model to incorporate some unknown quantity (external states beyond the blanket) that the system can do inferences about at all. Variational free energy, being a tractable (i.e., easily computed) upper bound on surprisal, already makes this a useful tool to have—however we will go further and claim supposing the existence of a minimised variational free energy, with its interpretation as inferring a target density by approximating it with a variational density, has a use in interpreting models in statistical physics in a new way.
We will remark that much of this story can be told for the trajectories of a system and distributions over path spaces, but with many modifications [11, 4], none of which are necessary to make the point sought in this paper.
@@ 2.2 Analysis and interpretation
Here, variational free energy is a Lyapunov function—a function that decreases monotonically along the trajectories of a (deterministic) dynamical system. The utility of Lyapunov functions is that they guarantee stability, allowing us to understand the dynamical system as flowing on the gradient of the Lyapunov function. Variational free energy arises as a Lyapunov function for the straightforward reason that the most probable state is the least surprising one, and that the log-probability is able to describe the probability of all fluctuations of the system in terms of their distance from that most likely state [22, 17, 23]—combined with the inequality in (2.1.2). Since the dynamics towards the mode induce a gradient flow on a divergence between the estimated parametric distribution and the true distribution, the FEP moreover proposes that these dynamics (in virtue of the coupling) can be modelled as a variational inference process, following a gradient flow on the variational free energy. Said another way, the crucial idea is that the dynamics of a system do inference when written as minimising a free energy functional, which is itself the consequence of reading surprisal as a Lyapunov function for those dynamics, following e.g. Freidlin–Wentzell, and noting that the KL divergence is stipulated to be minimised by a near-tautology that the system will be at its average state on average. In that case, we can map these parameters to $\sigma(\hat{\mu}_{b})$, meaning that for as long as the particle is at state $\hat{\mu}_{b}$, the particle models its environment by estimating the parameters of its distribution.
@@
Ultimately this simply means that, in virtue of being coupled (via blanket states) to its environment, the particle reflects data about its environment—or, what it believes that its environment is like. Under the assumption that a particle which exists in an environment models its environment, free energy is implicitly a measurement of coherence. If the particle stops modelling its environment in a particular way (i.e., for a particular $\sigma$), it must have ceased to exist with the mean that $\sigma$ was constructed for; conversely, if a particle ceases to exist, it will stop modelling its environment. The surprisal being considered is, much like probability, surprisal given a model, in that it quantifies how surprising it would be for an observer—equipped with a model—to find the system (such as itself) in some non-system-like state or configuration. Minimising surprisal implies that the variational free energy is minimised, (i.e., the object-as-model is a good estimator of its environment [11]), and likewise, minimising the variational free energy places an upper bound on the surprisal.
@@
Similarly, note the direction of the implication in our claims. In the context of modelling persistent systems, the FEP is a sort of inversion of the typical reasoning one sees: instead of saying an object persists (as a thing) because it is stable, we say an object is stable (as what it is) because it persists (as that thing) [16]. We stipulate that the KL divergence vanishes—because it is a tautology that $p$ should equal $q$ almost surely—and because physics only ‘works’ in a given regime if this is the case (see [11, §IIB] where this argument is spelt out). This happens in the background, so to speak. Then we can use the FEP to describe the dynamics as a gradient flow on surprisal—which can be done by a simple mathematical fact and independently of the FEP—as a process which has performed inference in virtue of existing. Adding this ‘inflationary’ KL term, the non-equilibrium steady state density of the system can be understood as a model of the environment under this stipulation. All this is made possible by the existence of a particular partition. Hence, we arrive at a deflationary account of intelligence through an inflationary mathematical technique. However—note that when the Markov blanket consists of sensory states and active states, this inference can immediately be written as part of a control problem [27, 28, 29, 30, 31, 32, 33], lending it an even more sophisticated interpretation.
@@ 3.1 Cellular morphogenesis
Starting from an ensemble of undifferentiated cells at the same location, we wanted to simulate their migration and differentiation using the minimisation of variational free energy. The simulation here exploits the fact one can express a Markov blanketed system’s marginal Lyapunov function in terms of a variational free energy function, such that its attracting set can be prescribed in terms of the generative model that defines the free energy. In essence, the systems dynamics rest upon each cell inferring its unique identity (in relation to all others) and behaving in accord with those inferences. [^1] These simulations were based on [34] and use the DEM toolbox in the MATLAB software SPM12 (https://www.fil.ion.ucl.ac.uk/spm/). They can be reproduced using the morphogenesis feature and the data provided there.
@@
The model was operationalised by following the procedure in [14]. A Taylor expansion of trajectories was taken and the statistics of each order of approximation $(\psi,\psi^{\prime},\psi^{\prime\prime},\ldots)$ were modelled at each point in time. A derivative operator $D$ was defined for ordinary differential equations on this space. The entire trajectory was given a coordinate in function space
@@
for each cell. [^2] This follows from the marginal flow lemma of [35]. This is where the interpretation as inference which we supply as modellers comes into play. We cannot compute the solution to (3.1.1) since the surprisal term yields an intractable integral. However, under the generative model we have equipped our system with, we can instead treat the dynamics of the cell under (3.1.1) as maximising log-evidence for the sensations of the cell, and hence as a variational problem—as posed throughout §2. We will introduce the variational distribution
@@
Treating (3.1.1) as a Bayesian filter of trajectories, from the derivation outlined in [14, equation 3.6] the final drift terms for the equations of motion for the $i$-th cell given the generative model were obtained as (again denoting the internal state $\psi_{i}$ as $\mu$)

### v5

@@ Acknowledgements
This paper more fully develops the theory outlined in a previous manuscript [mtff], and we are grateful to those listed there for their contributions to preliminary discussions on the topic. The account here is largely based on a series of talks given by the third named author at the Department of Philosophy of the University of Vienna, the Laufer Center for Physical and Quantitative Biology, the 2024 Conference on Complex Systems, and the informal discussions that followed. We thank the organisers and audiences of both. DARS acknowledges support from the Einstein Chair programme at the Graduate Centre of the City University of New York. KJF and DARS acknowledge support from the VERSES Research Lab.
@@ 1 Introduction
In statistics, generative models are joint probability distributions that model the relationships between different variables—most often, observations and their causes, or data and a set of labels. In the study of open and interacting systems, the use of statistical models—capable of describing such relationships—seems to be particularly apt. One approach to making such models, using variational (approximate Bayesian) inference [Friston2019, ramstead2022bayesian, friston2023path], describes how a physical object reflects the properties of its environment and vice versa. In this paper, we argue that generative models provide an especially useful foundation in scientific modelling generally, and especially in the study of systems exhibiting adaptation, morphogenesis, and other non-equilibrium phenomena. More specifically, we argue that generative modelling—premised on the use of variational inference to model the causes of the signals that a coupled system is receiving [beal2003variational]—is particularly apt for the practice of modelling the activity of modelling itself; for instance, in modelling sophisticated self-organising systems, where we write down a representation of the model that we assume is in play in prediction and action. We use the ensuing variational free energy principle to illustrate how implementing models which (i) formalise how one object keeps track of another, and (ii) codify these relationships, can increase the fidelity of our mathematical modelling.
@@
One of the points this paper will emphasise is that using this framework does not necessarily presuppose that a system actually performs inference on its environment. It is a truism that physical systems need not explicitly calculate their trajectories of motion, to be modelled as pursuing such trajectories. We need not assume that the particle itself is literally performing inference when we build models under our framework. What is at stake is an ‘as if’ description: we simply assume that there exists a quantity that varies systematically with the dynamics of the system; it so happens that surprisal (or, equivalently, variational free energy) is a good choice of quantity for physical reasons, and that the minimisation of this quantity is mathematically equivalent to inference, in the sense of being an estimator. The fact that it is our model of the system which does inference and not necessarily the system itself is both a boon and a bane. It enables us to model the system using the more tractable gradient of variational free energy and the mathematics of statistical inference, rather than studying the possibly highly non-linear coupling between the dynamics of the two systems. However, it is not necessarily a faithful representation of the system in the literalist sense. (See also [Andrews2021].)
@@
The structure of the paper is as follows. We first present a conceptual introduction to the FEP—and the way it leverages the relationships encoded by a generative model to build a scientific model of the dynamics of physical objects using free energy gradients. This is possible by modelling the property of physical systems that they look as if they engage in modelling to an external observer. To sketch the argument: the FEP provides us with tools from dynamical systems and information theory, allowing us to model any coupled random dynamical systems as statistical estimators, i.e., as engaging in the statistical estimation of some quantity pertaining to the set of objects with which it interacts. Under the FEP, one object that is coupled to any other can thus be read as a representation of the coupled or estimated object, albeit of a deflationary, qualified sort (discussed in [Ramstead2020semantics]; see also [Ramstead2019enactive, ramstead2022bayesian]). More formally, the FEP applies to generative models that contain a partition of the states or paths of a joint system (a so-called ‘particular’ partition, i.e., into objects or particles). This partition induces a statistical boundary called a Markov blanket [Pearl1998], which captures conditional independence relations amongst subsets of the system [Friston2019, sakthivadivel2022d]. The FEP says that, when such Markov blanketed subsets of the system exist, those subsystems track each others’ statistical properties (dynamically) across the blanket [classical-physics]. Essentially, the FEP is a modelling framework which captures the following fact: models of coupled particles or factorised systems can be written as parameter estimators, and estimating the parameters of a distribution is a form of Bayesian inference.
@@
We then argue that there is no ambiguity between model and modelled real-world system in this formalism. We can think of the FEP, metaphorically, as a map of the part of the territory that behaves as if it were a map—or looks as such to an external observer. Accounts such as [Andrews2021] have encouraged practitioners of the FEP to be explicit about this point. Viewing the theory this way, not only is there no conflation between our FEP-theoretic model of the system and the system itself, but in fact we arrive at a novel modelling method for coupled random dynamical systems. In particular, we do not claim that the objects in the system themselves must literally be performing inference. Rather, we describe the dynamics as doing inference, in the sense of being a statistical estimator whose states parameterise a variational density.
@@
In summary we propose that the FEP provides a map of any possible map whatsoever—of, or held by—a physical system. (This includes the use of the FEP itself by a modeller.) We suggest that this leads to constraints on all possible maps, or models of physical things, arising from what it means to be a map or model of a physical process at all. In the fashion of Jaynes, who proposed that doing statistical mechanics is doing inference about the probabilities of microstates given knowledge of macrostates [jaynes1, jaynes2], the FEP is a natural expression of statistical physics as a process of making models. More broadly we claim that this echoes the sentiments of Wittgenstein, providing a constraint on modelling in terms of logically necessary preconditions for sense-making and the parsing of sensory streams [Ramstead2020semantics].
@@
For those already familiar with the FEP, this paper can be thought of as going back to basics, emphasising its foundation in inverting generative models (written as stochastic differential equations (SDEs) in generalised coordinates), using a path integral scheme for minimising variational free energy [friston2010generalised].
@@ 2.1 Fundamentals
Here we will describe the proposed free energy principle (also directing the reader to prior literature such as [classical-physics, ramstead2022bayesian, friston2023path] for a more extended look). The FEP aims to describe the physical properties of a system in terms of dynamics on a manifold of probability distributions (i.e., as variational inference) [DaCosta2021], resting on the useful property that physical objects generally look as if they track or infer the statistics of things to which they are coupled [ramstead2022bayesian, Sakthivadivel2022b, classical-physics]. Formulated most generally, the FEP says that the average path taken by an object that is sparsely coupled to its environment is characterised by a minimisation of a variational free energy functional.
We consider a particle (also referred to as a ‘system’) with states $\mu$ coupled to another system with states $\eta$ (henceforth, the ‘environment’) through a set of variables $B$ with values $b$, and a steady state density $p(\eta,b,\mu)$ such that $\mu$ is conditionally independent of $\eta$ given $B$. After Pearl (see [Pearl1998]) we refer to $B$ as the Markov blanket of the system. Typically, $B$ will itself consist of two different kinds of states—sensory states, providing the system signals about its environment, and active states, providing interventions of the system on the environment. The coupling at play puts us in a non-equilibrium situation, where a system can be driven by fluxes of energy or matter from other systems in its environment [seifert2012stochastic].
The FEP begins by saying that if there exists a boundary in the form of a Markov blanket, then there exists a function $\sigma$ relating (the dynamics of) internal states to the environment across the boundary, and in particular, such that $\sigma(\hat{\mu}_{b})=\hat{\eta}_{b}$, where $\hat{\,\cdot\,}_{b}$ denotes the conditional expectation given the blanket state $b$ [Sakthivadivel2022b, Lemma 4.3]. Following that, the aim of the FEP is to understand what the generative model $p(\eta,b,\mu)$ says about the dynamics of the subsystem with states $\mu$, beyond just its marginal distribution. In particular: if we assume that $\hat{\eta}_{b}$ parametrises $p(\eta\mid b)$, then we have made $\mu$ into an estimator of the statistics of the environment.
@@
in terms of a gradient flow on the surprisal weighted by a particular matrix field [ao2004potential, ma2015, yuan2017sde, da2023entropy], having the form
@@
Moreover, the log-probability is able to describe the probability of all fluctuations of the system in terms of their distance from the most likely state, as well as physically-derived quantities [freidlin_wentzell_2012, seifert2012stochastic, seifert2019stochastic], making this a meaningful equation. (See [parr2020markov] where this is applied to reproduce certain aspects of stochastic thermodynamics.)
@@
Inference by parameter estimation implies the Kullback–Leibler (KL) divergence between the posterior and the true density vanishes as the parameter varies. We will write
@@
with equality precisely when the KL term vanishes. We conclude that the minimisation of surprisal implies the vanishing of the divergence term. Let us apply the Laplace approximation [friston2007variational] such that the expected state is approximately the mode of the density. It follows that the most probable state is the least surprising one. Hence, for physical reasons, we expect the surprisal to be minimised. From that reasoning, (2.1.1) has the equivalent form
@@
What we have done is used the fact that the dynamics of a (steady state) system minimise surprisal on average, and argued that whenever equipped with a Markov blanket they also minimise variational free energy; and hence, the dynamics of a system with a particular partition can always be written as a gradient flow on variational free energy. This gradient flow is what we identify as actually being the process (i.e., the performance) of inference. Explicit examples of this can be found in [parr2020markov, ueltzhoffer2021drive, classical-physics]. The particular partition becomes crucial here, since this is what we leverage in our model to incorporate some unknown quantity (external states beyond the blanket) that the system can do inferences about at all. Variational free energy, being a tractable (i.e., easily computed) upper bound on surprisal, already makes this a useful tool to have—however we will go further and claim supposing the existence of a minimised variational free energy, with its interpretation as inferring a target density by approximating it with a variational density, has a use in interpreting models in statistical physics in a new way.
We will remark that much of this story can be told for the trajectories of a system and distributions over path spaces, but with many modifications [classical-physics, friston2023path], none of which are necessary to make the point sought in this paper.
@@ 2.2 Analysis and interpretation
Here, variational free energy is a Lyapunov function—a function that decreases monotonically along the trajectories of a (deterministic) dynamical system. The utility of Lyapunov functions is that they guarantee stability, allowing us to understand the dynamical system as flowing on the gradient of the Lyapunov function. Variational free energy arises as a Lyapunov function for the straightforward reason that the most probable state is the least surprising one, and that the log-probability is able to describe the probability of all fluctuations of the system in terms of their distance from that most likely state [freidlin_wentzell_2012, seifert2012stochastic, seifert2019stochastic]—combined with the inequality in (2.1.2). Since the dynamics towards the mode induce a gradient flow on a divergence between the estimated parametric distribution and the true distribution, the FEP moreover proposes that these dynamics (in virtue of the coupling) can be modelled as a variational inference process, following a gradient flow on the variational free energy. Said another way, the crucial idea is that the dynamics of a system do inference when written as minimising a free energy functional, which is itself the consequence of reading surprisal as a Lyapunov function for those dynamics, following e.g. Freidlin–Wentzell, and noting that the KL divergence is stipulated to be minimised by a near-tautology that the system will be at its average state on average. In that case, we can map these parameters to $\sigma(\hat{\mu}_{b})$, meaning that for as long as the particle is at state $\hat{\mu}_{b}$, the particle models its environment by estimating the parameters of its distribution.
@@
Ultimately this simply means that, in virtue of being coupled (via blanket states) to its environment, the particle reflects data about its environment—or, what it believes that its environment is like. Under the assumption that a particle which exists in an environment models its environment, free energy is implicitly a measurement of coherence. If the particle stops modelling its environment in a particular way (i.e., for a particular $\sigma$), it must have ceased to exist with the mean that $\sigma$ was constructed for; conversely, if a particle ceases to exist, it will stop modelling its environment. The surprisal being considered is, much like probability, surprisal given a model, in that it quantifies how surprising it would be for an observer—equipped with a model—to find the system (such as itself) in some non-system-like state or configuration. Minimising surprisal implies that the variational free energy is minimised, (i.e., the object-as-model is a good estimator of its environment [classical-physics]), and likewise, minimising the variational free energy places an upper bound on the surprisal.
@@
Similarly, note the direction of the implication in our claims. In the context of modelling persistent systems, the FEP is a sort of inversion of the typical reasoning one sees: instead of saying an object persists (as a thing) because it is stable, we say an object is stable (as what it is) because it persists (as that thing) [Sakthivadivel2022b]. We stipulate that the KL divergence vanishes—because it is a tautology that $p$ should equal $q$ almost surely—and because physics only ‘works’ in a given regime if this is the case (see [classical-physics, §IIB] where this argument is spelt out). This happens in the background, so to speak. Then we can use the FEP to describe the dynamics as a gradient flow on surprisal—which can be done by a simple mathematical fact and independently of the FEP—as a process which has performed inference in virtue of existing. Adding this ‘inflationary’ KL term, the non-equilibrium steady state density of the system can be understood as a model of the environment under this stipulation. All this is made possible by the existence of a particular partition. Hence, we arrive at a deflationary account of intelligence through an inflationary mathematical technique. However—note that when the Markov blanket consists of sensory states and active states, this inference can immediately be written as part of a control problem [friston2012active, pezzulo2015active, pio2016active, pezzulo2018hierarchical, baltieri2019pid, millidge2020relationship, baioumy2021active], lending it an even more sophisticated interpretation.
@@ 3.1 Cellular morphogenesis
Starting from an ensemble of undifferentiated cells at the same location, we wanted to simulate their migration and differentiation using the minimisation of variational free energy. The simulation here exploits the fact one can express a Markov blanketed system’s marginal Lyapunov function in terms of a variational free energy function, such that its attracting set can be prescribed in terms of the generative model that defines the free energy. In essence, the system’s dynamics rest upon each cell inferring its unique identity (in relation to all others) and behaving in accord with those inferences. [^1] These simulations were based on [friston2015knowing] and use the DEM toolbox in the MATLAB software SPM12 (https://www.fil.ion.ucl.ac.uk/spm/). They can be reproduced using the morphogenesis feature and the data provided there.
@@
The model was operationalised by following the procedure in [friston2010generalised]. A Taylor expansion of trajectories was taken and the statistics of each order of approximation $(\psi,\psi^{\prime},\psi^{\prime\prime},\ldots)$ were modelled at each point in time. A derivative operator $D$ was defined for ordinary differential equations on this space. The entire trajectory was given a coordinate in function space
@@
for each cell. [^2] This follows from the marginal flow lemma of [Friston2019]. This is where the interpretation as inference which we supply as modellers comes into play. We cannot compute the solution to (3.1.1) since the surprisal term yields an intractable integral. However, under the generative model we have equipped our system with, we can instead treat the dynamics of the cell under (3.1.1) as maximising log-evidence for the sensations of the cell, and hence as a variational problem—as posed throughout §2. We will introduce the variational distribution
@@
Treating (3.1.1) as a Bayesian filter of trajectories, from the derivation outlined in [friston2010generalised, equation 3.6] the final drift terms for the equations of motion for the $i$-th cell given the generative model were obtained as (again denoting the internal state $\psi_{i}$ as $\mu$)